1. The United States can produce one DVD player in 4 hours and one bushel of corn in 10 hours. Canada, on the other hand, produces one DVD player in 5 hours and one bushel of corn in 20 hours. According to the Law of Comparative Advantages, the United States would benefit by specializing in:

a. DVD players

b. Corn

c. none

2. The United States can produce one cell phone in 10 hours and one barrel of ethanol fuel in 15 hours. Canada, on the other hand, produces one cell phone in 12 hours and one barrel of ethanol fuel in 25 hours. According to the Law of Comparative Advantages, Canada would benefit by specializing in:

a. cell phones

b. ethanol

c. none
Answer #1

1) In US

OC of DVD player = 10/4=2.5

OC of corn = 4/10 = 0.4

In Canada

OC of DVD player = 20/5 = 4

OC of corn = 5/20 = 0.25

The United States would benefit by specializing in:

a. DVD players

2) In US

OC of Cell phone = 15/10 = 1.5

OC of fuel = 10/15 = 0.6

In Canada

OC of cell phone = 25/12 = 2

OC of fuel = 12/25 = 0.48

Canada would benefit by specializing in:

b. ethanol fuel
Answer #2

Answers Based on Comparative Advantage:


Question 1: U.S. Specialization (DVD Players vs. Corn)

  • U.S. Opportunity Costs:

    • 1 DVD player: 4 hrs10 hrs (corn)=0.4 bushels.

    • 1 bushel corn: 10 hrs4 hrs (DVD)=2.5 DVD players.

  • Canada’s Opportunity Costs:

    • 1 DVD player: 5 hrs20 hrs (corn)=0.25 bushels.

    • 1 bushel corn: 20 hrs5 hrs (DVD)=4 DVD players.

  • Comparative Advantage:

    • The U.S. has a lower opportunity cost in DVD players (0.4 bushels vs. Canada’s 0.25).

    • Canada has a lower opportunity cost in corn (4 DVDs vs. U.S.’s 2.5).


  • Answer:
    The U.S. should specialize in a. DVD players.



Question 2: Canada’s Specialization (Cell Phones vs. Ethanol)

  • Canada’s Opportunity Costs:

    • 1 cell phone: 12 hrs25 hrs (ethanol)=0.48 barrels.

    • 1 barrel ethanol: 25 hrs12 hrs (cell)2.08 cell phones.


  • U.S. Opportunity Costs:

    • 1 cell phone: 10 hrs15 hrs (ethanol)0.67 barrels.

    • 1 barrel ethanol: 15 hrs10 hrs (cell)=1.5 cell phones.


  • Comparative Advantage:

    • Canada has a lower opportunity cost in ethanol (2.08 cell phones vs. U.S.’s 1.5).

    • The U.S. has a lower opportunity cost in cell phones (0.67 barrels vs. Canada’s 0.48).


  • Answer:
    Canada should specialize in b. ethanol.


answered by: anonymous